# KiDS-Legacy: Redshift distributions and their calibration

### v1

@@ 4.3 Prior redshift weight
The SOM implementation of the direct calibration method is designed to perform two primary tasks: reweight the colour space of the calibration sample to better represent the wide-field sample, and flagging wide-field sources for removal where this correction is not possible (i.e. the gold-weighting).
These corrections assume, however, that the probability distribution of redshift at a given colour in the calibration and wide-field samples are identical. Such an assumption is easily violated in the process of spectroscopic redshift acquisition, where two galaxies with different redshift but the same broadband colours (i.e. those with colour-redshift degeneracy) have different spectroscopic redshift success rates (as, e.g., one galaxy shows the [OII] doublet in the optical and the other does not). Such a selection in the successful acquisition of spectroscopic redshifts has been shown to lead to pathological biases in vanilla direct calibration implementations (Hartley et al., 2020).
Even more simply, however, this assumption is also easily violated when the samples are constructed from vastly different selection functions (e.g. Gruen & Brimioulle, 2017). For example, two simple magnitude-limited samples constructed from different magnitude limits will probe different redshift baselines. If the deeper sample has access to galaxies which are colour-degenerate with galaxies in the shallower window, then the redshift distribution at fixed colour will be unimodal for the shallow sample and multimodal for the deeper sample.
Correcting for this effect is complicated, as it requires one to know the distribution of redshift for the target sample of galaxies (which is our desired end-product of the SOM calibration process). In KiDS-Legacy, we perform a first order correction using an a priori estimate of the wide-field sample redshift distributions (see below), and remove significant differences between this wide-field estimate and the (known) redshift distribution of the calibration sample.
To perform this correction, we require an estimate of the true redshift distribution of the cosmic shear wide-field galaxy sample. To this end we start by constructing an analytic expression for the redshift distribution of an arbitrary magnitude limited sample. Using the raw $108$ $\text{deg}^{2}$ SURFS-Shark lightcone (see Sect. 3), which contains noiseless SDSS and VIKING fluxes, we construct samples of galaxies cut to various magnitude limits in a range of photometric bands. We then fit each of the resulting galaxy samples with the function:
@@
where $A$ and $\alpha$ are free parameters. We then model $A(m)$ and $\alpha(m)$ with a fourth-order polynomial.
@@
This allows us to construct an analytic redshift distribution for a sample of galaxies that is magnitude limited (in true flux) between 18th and 27th magnitude in any band from $u$ to $Z$. An example showing the estimated model parameters, the polynomial fits, and the resulting analytic $N(z)$ is given in Fig. 5.
@@
Figure 5: Model parameters and the resulting analytic $N(z)$ for samples defined as magnitude limited (in the $r$-band) from our noiseless SURFS+Shark lightcone. Panels left and centre show the free parameters from Eq. (2), as a function of the $r$-band magnitude limit, including polynomial fits. The right panel shows the analytically estimated $N(z)$ for each of the models parameters in the other two panels.
@@
We subsequently construct prior volume corrective weights for the KiDS-Legacy calibration sample by first producing an analytic approximation to the wide-field sample redshift distribution using our analytic prescription, using a magnitude limited sample that we believe most closely mimics the true selection function of the wide-field data. This is complicated by the various complex lensing selections (and shape weights) that are applied to the wide field lensing sample of galaxies.
@@
We choose to use a sample that is magnitude limited in the $r$-band, at $20\leq r\leq 23.5$. The bright-end magnitude limit is chosen because of the selection performed by lensfit  to limit galaxies to those with $r\geq 20$. The faint-end limit is chosen due to the lensing weights returned by lensfit  which are strongly magnitude dependent: at $r\approx 23.5$ the lensing weight is roughly half its maximum. We define the corrective prior volume weights as the ratio of the redshift distribution PDFs $P_{w}(z)/P_{c}(z)$, where $w$ and $c$ refer to the analytic wide-field sample and the data calibration sample respectively. The impact of the prior volume weights on the total $N(z)$ of the spectroscopic compilation, and also for an example tomographic bin, are shown in Fig. 6. The figures show the distributions before SOM weighting. It is clear from the figure that the prior volume weights have a systematic effect on the relative weight of individual calibrating sources as a function of redshift and, perhaps more importantly, that this manifests as a shift in the entire pre-weighting $N(z)$ for some tomographic bins.
Figure 6: Impact of prior-volume weights on the (pre-SOM) $N(z)$ of the full spectroscopic compilation (upper panel) and on those galaxies in the compilation that end up in tomographic bin three (lower panel). The systematic effect that the prior weight imparts on the tomographic bin is particularly clear, where the shift in probability mass from $z\approx 0.9$ to $z\approx 0.5$ creates a systematic shift in the full tomographic bin three $N(z)$.
@@ 4.4 Redshift distribution bias estimation
Calibration of the redshift estimation process (specifically the derivation of tomographic bin redshift bias parameters) in KiDS-Legacy is performed by implementing the redshift distribution estimation pipeline on our various simulated data sets, which are designed to mimic the observed data as accurately as possible. With the estimated redshift distributions, and the known true (weighted) redshift distribution of the source samples, we then compute the bias of each estimated redshift distribution as:
@@
where $\mu_{z}$ is the (shape- and gold-) weighted mean true redshift of the wide-field sample, $\hat{\mu}_{z}$ is the estimated mean redshift of the wide-field sample, computed directly from our weighted calibration sample [^8] For our fiducial $N(z)$ binning, $\Delta z=0.05$, the primary probability mass of each $N(z)$ is sampled by ten or more bins. This means that the difference introduced when computing the sample mean redshift vs $N(z)$ expectation is negligible.. We typically perform this measurement using many realisations of the calibration samples, which produces many estimates of $\hat{\mu}_{z}$ (and, because of the gold selection/weighting, possibly many different $\mu_{z}$). Our final quoted biases are the arithmetic means of the biases estimated per tomographic bin and/or pipeline setup ($\langle\delta z\rangle$). The population scatter of the biases in these realisations is also a relevant consideration, and is quoted as $\sigma_{\delta z}$. We note in particular that these uncertainties are smaller than those in previous KiDS analyses, due to the change in simulation philosophy described in Sect. 3.3.
@@ 5 Clustering redshift methodology
As a complementary approach to test and validate the SOM $N(z)$  we use clustering redshifts (e.g. Newman, 2008) following previous KiDS work (Hildebrandt et al., 2017, 2020, 2021; Morrison et al., 2017; van den Busch et al., 2020). The dr5 analysis presented here is an evolution of these previous works adding more area for the measurements of the cross-correlations and – more importantly – expanding the suite of external spectroscopic surveys used for the calibration. The dedicated KiDZ data allow us to measure cross-correlations with VIPERS, providing additional constraints in the range $0.6<z\lesssim 1.2$. Recently, the DESI Early Data Release provided a lot of additional galaxies with spectroscopic redshifts extending to even higher redshifts and overlapping with the KiDS main survey area in six rosette-shaped DESI pointings. Together these advances allow us to validate the SOM $N(z)$ for the first five tomographic bins (bin six is only partly covered) without referring to deep, pencil-beam surveys (as we still had to do in Hildebrandt et al., 2021), significantly decoupling the clustering redshift from the SOM approach in terms of calibration data. While VIPERS is also used for the SOM calibration, its contribution to both calibrations is small (see Sect. 5.2).
@@ 5.1 Correlation measurements
KiDS clustering redshifts are estimated with the versatile, public code yet_another_wizz [^9] https://pypi.org/project/yet-another-wizz/ (YAW; van den Busch et al., 2020), which is based on concepts already introduced by Schmidt et al. (2013) and Morrison et al. (2017). In particular, we use the publicly available version 2.6.0 that differs from the versions used in previous publications (e.g. Hildebrandt et al., 2021; Naidoo et al., 2023) in a few ways.
The code now generates spatial regions per calibration sample based on $k$-means clustering of sky coordinates (e.g. using random catalogues) instead of splitting the data into individual pointings. These regions are used to estimate the data covariance via a spatial jackknife (previously using bootstrap). This empirical data covariance was tested against analytical models for the covariance matrix based on halo occupation distributions and a halo model approach for a different calibration data set derived from MICE2. While there is good general agreement between the features of the empirical and the analytical covariance, we opt to rely on the jackknife method due to the highly non-linear regime of the clustering measurements and possible uncertainties in the halo occupation distribution as well as non-Limber effects in the connected non-Gaussian terms (for details we refer to section 9.2 of Reischke et al., 2024). Regardless, the agreement with the analytic covariance serves as a good cross-check for the empirical jackknife covariance that we use throughout the clustering redshift analysis.
In addition to that, the code is now measuring pair counts across the boundaries of these spatial regions [^10] While counting pairs only within the same region can have an effect on the overall correlation amplitude, it has no effect on any of our previously published results because this overall amplitude is later normalised. and the Landy-Szalay estimator (Landy & Szalay, 1993) is used for all auto-correlation measurements. Finally, in the case of the cross-correlations, which use the Davis & Peebles (1983) estimator, only one random catalogue is needed. Hence, one can decide whether to use a random catalogue for the spectroscopic or KiDS data. We decide to use random catalogues for the spectroscopic data instead of the KiDS data in those cases, since most of the spectroscopic surveys provide well-established random catalogues that take into account and correct for a lot of systematic effects.
@@ 5.2 Fiducial analysis setup
We measure angular correlations in a single bin of fixed transverse physical separation between $0.5<r\leq 1.5\leavevmode\nobreak\ \mathrm{Mpc}$ in 32 linearly spaced redshift bins in the range $1.60.05\leq z<$ 1.6.111111 These scales are larger than what was used in previous KiDS clustering redshift analyses. See Sect. 6.2.2 for a justification. For MICE2, the upper redshift limit is reduced to $z_{\rm max}=1.4$. Furthermore, we need to mitigate the redshift evolution of the galaxy bias of the calibration data and the KiDS-Legacy data set. Following the notation of van den Busch et al. (2020), these biases can – under certain assumptions – be expressed in terms of the amplitudes of the angular autocorrelation functions of our spectroscopic reference sample, $w_{\rm ss}(z)$, and our wide-field photometric sample, $w_{\rm pp}(z)$. Then, the true (unknown) redshift distribution can be written as
@@
where $w_{\rm sp}(z)$ is the cross-correlation amplitude between our spectroscopic and photometric samples, $\Delta z$ is the bin width of the CC measurements, and $N_{\rm CC}(z)$ denotes our estimated $N(z)$ from the cross-correlation method. In our fiducial analysis we choose to only correct for the calibration data bias, i.e. we effectively measure the numerator of the right-hand side of Eq. (4). The impact of the unknown galaxy bias of the KiDS-Legacy data is partly mitigated by the fact that we bin the data tomographically (Schmidt et al., 2013) and it has been shown previously to be sufficiently small that we can neglect it in our following analysis (see e.g. Hildebrandt et al., 2021; van den Busch et al., 2020).
Finally, we must produce a joint redshift estimate from all five calibration samples for which we measure the correlation functions independently. We obtain this final estimate by computing the inverse-variance weighted average of the bias-corrected correlation measurements of all reference samples. This produces an optimal redshift estimate that reflects the redshift-dependent relative contribution of each correlation measurement to the overall redshift estimate and accounts for the different statistical power each calibration sample has at any given redshift (Fig. 7). In general, the joint CCs are dominated by BOSS, GAMA, and 2dFLenS at $z\lesssim 0.7$ and by DESI at high redshifts, whereas VIPERS has an overall small contribution due to its limited total area and number density.
Figure 7: Example of the inverse-variance weighted combination of CCs computed from the fourth tomographic bin of MICE2. The top panel shows the individual measurements from each calibration sample, the middle panel the weighted average, and the bottom panel the relative weight of each sample as a function of redshift.
@@ 5.3 Adaptation of signal-to-noise in MICE2
When comparing the signal-to-noise ratio in the measured CCs on the KiDS/KiDZ data and MICE2, we find that the noise level in the simulation is typically much smaller in most redshift bins, especially in the sixth tomographic bin (see top panel of Fig. 8). The reason for this difference is not entirely clear. Possible reasons could be differences in the selection functions applied to MICE2 for both, the calibration and KiDS-Legacy data, as compared to the real data. There may also be a difference in the overall clustering amplitude found in the data and the simulations, especially on small scales. Additionally, the key difference between the data and MICE2 is that the mock photometry is perfectly uniform such that the effects of variable depth (Heydenreich et al., 2020), are not present in the simulation. Finally, there may be other systematic effects in the data for which our MICE2 data does not account, e.g. related to fibre-collisions, which for example in DESI require special pair-weights (Bianchi et al., 2018) that we currently cannot integrate into our correlation estimator.
Since there is no clear explanation for this difference between data and simulation, we opt to adapt the measurements on the mocks by adding additional Gaussian noise to the values and inflating the error bars such that they match the data. We compute the right amount of noise required from the difference between the covariances of the inverse-variance-combined measurements of data and mock. Figure 8 shows a comparison of the original measurements in MICE2 and how the adapted version compares to the data measurements in the sixth tomographic bin.
Figure 8: One realisation of the CC measurements from MICE2 in the sixth tomographic bin after noise adaptation (Sect. 5.3) and adding intrinsic scatter (Sect 5.4). The top panel compares just the uncertainties of $N_{\rm CC}(z)$ from the data (grey) and MICE2 before (blue) and after adaptation (red). The black dashed line indicates the fitted intrinsic scatter from the measurements on the data ($f=0.14\pm 0.04$; see Eq. 6). The bottom panel shows the measured $N_{\rm CC}(z)$ from MICE2 (blue points) compared to the data (grey line). The red data points represent the MICE2 measurements after adapting the noise (upscaling errors and perturbing values) and adding the intrinsic scatter (only perturbing values) obtained from the data.

### v2

@@ 4.3 Prior redshift weight
The SOM implementation of the direct calibration method is designed to perform two primary tasks: re-weight the colour space of the calibration sample to better represent the wide-field sample, and flag wide-field sources for removal where this correction is not possible (i.e. the gold-weighting). These corrections assume, however, that the probability distribution of redshift at a given colour in the calibration and wide-field samples are identical. Such an assumption is easily violated in the process of spectroscopic redshift acquisition, where two galaxies with different redshift but the same broadband colours (i.e. those with colour-redshift degeneracy) have different spectroscopic redshift success rates (e.g. one galaxy shows the [OII] doublet in the optical and the other does not). Such a selection in the successful acquisition of spectroscopic redshifts has been shown to lead to pathological biases in vanilla direct calibration implementations (Hartley et al., 2020).
Even more simply, however, this assumption is also easily violated when the samples are constructed from vastly different selection functions (e.g. Gruen & Brimioulle, 2017). For example, two simple magnitude-limited samples constructed from different magnitude limits will probe different redshift baselines. If the deeper sample has access to galaxies that are colour-degenerate with galaxies in the shallower window, then the redshift distribution at fixed colour will be unimodal for the shallow sample and multi-modal for the deeper sample.
Correcting for this effect is complicated, as it requires one to know the distribution of redshift for the target sample of galaxies (which is our desired end-product of the SOM calibration process). In KiDS-Legacy, we performed a first order correction using an a priori estimate of the wide-field sample redshift distributions (see below) and removed significant differences between this wide-field estimate and the (known) redshift distribution of the calibration sample.
To perform this correction, we required an estimate of the true redshift distribution of the cosmic shear wide-field galaxy sample. To this end we began by constructing an analytic expression for the redshift distribution of an arbitrary magnitude-limited sample. Using the raw $108$ $\text{deg}^{2}$ SURFS-Shark light cone (see Sect. 3), which contains noiseless SDSS and VIKING fluxes, we constructed samples of galaxies cut to various magnitude limits in a range of photometric bands. We then fitted each of the resulting galaxy samples with the function:
@@
where $A$ and $\alpha$ are free parameters. We then modelled $A(m)$ and $\alpha(m)$ with a fourth-order polynomial. This allowed us to construct an analytic redshift distribution for a sample of galaxies that is magnitude-limited (in true flux) between the 18th and 27th magnitude in any band from $u$ to $Z$. An example showing the estimated model parameters, the polynomial fits, and the resulting analytic $N(z)$ is given in Fig. 5.
@@
Figure 5: Model parameters and the resulting analytic $N(z)$ for samples defined as magnitude-limited (in the $r$ band) from our noiseless SURFS+Shark light cone. The left and centre panels show the free parameters from Eq. (2), as a function of the $r$-band magnitude limit, including polynomial fits. The right panel shows the analytically estimated $N(z)$ for each of the models parameters in the other two panels.
@@
We subsequently constructed prior volume corrective weights for the KiDS-Legacy calibration sample using this analytic prescription. To do this we first selected a magnitude-limited sample that we found closely mimics (in terms of the redshift distribution) the expected true selection function of our wide-field shear dataset; a process that is complicated by the various complex lensing selections (and shape weights) that are applied to the wide field lensing sample of galaxies.
@@
We chose to use a sample that is magnitude-limited in the $r$ band, at $20\leq r\leq 23.5$. The bright-end magnitude limit was chosen because of the selection performed by lensfitthat rejects galaxies at $r<20$. The faint-end limit was chosen due to the lensing weights returned by lensfit, which are strongly magnitude-dependent: at $r\approx 23.5$ the lensing weight is roughly half its maximum. With our analytic estimate of the wide-field sample redshift distribution, we then defined the corrective prior volume weights as the ratio of the redshift distribution PDFs $P_{w}(z)/P_{c}(z)$, where $w$ and $c$ refer to the analytic wide-field sample and the data calibration sample, respectively. The impact of the prior volume weights on the total $N(z)$ of the spectroscopic compilation, and also for an example tomographic bin, are shown in Fig. 6. The figures show the distributions both before and after SOM weighting, and when including our simulation-informed estimates of $N(z)$ bias (see Sect. 6). It is clear from the figure that the prior volume weights have a systematic effect on the relative weight of individual calibrating sources as a function of redshift and, perhaps more importantly, that this manifests as a shift in the entire pre-weighting $N(z)$ for some tomographic bins. After weighting by the SOM, however, the differences are less pronounced, indicating that some of the difference between the redshift distributions is absorbed by the colour-cell DIR weights (thanks to the colour-redshift relation). Finally, after incorporation of our simulation $N(z)$ bias estimates the differences are reduced even further (see Sects. 4.4 and 6), demonstrating the importance of having realistic simulations to correct for methodological bias in $N(z)$ estimation and calibration.
Figure 6: Impact of prior-volume weights on the (pre-SOM) $N(z)$ of the full spectroscopic compilation (upper panel), and on those galaxies in the compilation that end up in tomographic bin three both prior to the SOM re-weighting (middle panel) and after the SOM re-weighting (bottom panel). The systematic effect that the prior weight imparts on the tomographic bin is particularly clear prior to re-weighting by the SOM to match the wide-field sample but is somewhat diluted by the SOM re-weighting process. Nonetheless, there is a residual difference after the SOM re-weighting in the bottom panel. This difference is exacerbated if one does not have high quality image simulations to calibrate the underlying $N(z)$ bias (dashed lines in the bottom figure; discussed further in Sect. 6), as the non-prior weighted $N(z)$ estimate contains a significant systematic bias ($\langle\delta z\rangle\approx 0.03$), whereas the prior weighted $N(z)$ estimate has no bias ($\langle\delta z\rangle\approx 0.004$).
@@ 4.4 Redshift distribution bias estimation
Calibration of the redshift estimation process (specifically the derivation of tomographic bin redshift bias parameters) in KiDS-Legacy was performed by implementing the redshift distribution estimation pipeline on our various simulated datasets, which are designed to mimic the observed data as accurately as possible. With the estimated redshift distributions and the known true (weighted) redshift distribution of the source samples, we then computed the bias of each estimated redshift distribution as
@@
where $\mu_{z}$ is the (shape- and gold-) weighted mean true redshift of the wide-field sample, and $\hat{\mu}_{z}$ is the estimated mean redshift of the wide-field sample, computed directly from our weighted calibration sample [^8] For our fiducial $N(z)$ binning, $\Delta z=0.05$, the primary probability mass of each $N(z)$ is sampled by ten or more bins. This means that the difference introduced when computing the sample mean redshift vs $N(z)$ expectation is negligible.. We typically perform this measurement using many realisations of the calibration samples, which produces many estimates of $\hat{\mu}_{z}$ (and, because of the gold selection or weighting, possibly many different $\mu_{z}$). Our final quoted biases are the arithmetic means of the biases estimated per tomographic bin and pipeline setup ($\langle\delta z\rangle$). The population scatter of the biases in these realisations is also a relevant consideration, and is quoted as $\sigma_{\delta z}$. We note in particular that these uncertainties are smaller than those in previous KiDS analyses, due to the change in simulation philosophy described in Sect. 3.3.
@@ 5 Clustering redshift methodology
As a complementary approach to test and validate the SOM $N(z)$  we used clustering redshifts (e.g. Newman, 2008) following previous KiDS work (Hildebrandt et al., 2017, 2020, 2021; Morrison et al., 2017; van den Busch et al., 2020). The dr5 analysis presented here is an evolution of these previous studies, adding more area for the measurements of the CCs and, more importantly, expanding the suite of external spectroscopic surveys used for the calibration. The dedicated KiDZ data allowed us to measure CCs with VIPERS, providing additional constraints in the range $0.6<z\lesssim 1.2$. Recently, the DESI Early Data Release provided a number of additional galaxies with spectroscopic redshift estimates extending to even higher redshifts and overlapping with the KiDS main survey area in six rosette-shaped DESI pointings. Together these advances allowed us to validate the SOM $N(z)$ for the first five tomographic bins (bin six is only partly covered) without referring to deep, pencil-beam surveys (as we still had to do in Hildebrandt et al., 2021), significantly decoupling the clustering redshift from the SOM approach in terms of calibration data. One caveat, though, is VIPERS, which is used for both CCs and SOM $N(z)$ calibration. However we found that the overall contribution of VIPERS both SOM and CC $N(z)$ calibrations is subdominant (see Sect. 5.2) and so the independence of the methods is maintained.
@@ 5.1 Correlation measurements
KiDS clustering redshifts are estimated with the versatile, public code yet_another_wizz [^9] https://pypi.org/project/yet-another-wizz/ (YAW, van den Busch et al., 2020), which is based on concepts already introduced by Schmidt et al. (2013) and Morrison et al. (2017). In particular we used the publicly available version 2.6.0, which differs from the versions used in previous publications (e.g. Hildebrandt et al., 2021; Naidoo et al., 2023) in a few ways.
The code now generates $N$ spatial regions per calibration sample based on $k$-means clustering of sky coordinates (e.g. using random catalogues) instead of splitting the data into the $N$ individual square-degree pointings spanned by each dataset. These regions are used to estimate the data covariance via a spatial jackknife (previously using bootstrap). This empirical data covariance was tested against analytical models for the covariance matrix based on halo occupation distributions and a halo model approach for a different calibration dataset derived from MICE2. While there is good general agreement between the features of the empirical and the analytical covariance, we opt to rely on the jackknife method due to the highly non-linear regime of the clustering measurements, possible uncertainties in the halo occupation distribution, and non-Limber effects in the connected non-Gaussian terms (for details we refer to Section 9.2 of Reischke et al., 2025). Regardless, the agreement with the analytic covariance serves as a good cross-check for the empirical jackknife covariance that we use throughout the clustering redshift analysis.
In addition to the above, the code now measures pair counts across the boundaries of these spatial regions [^10] While counting pairs only within the same region can have an effect on the overall correlation amplitude, it has no effect on any of our previously published results because this overall amplitude is later normalised. and the Landy-Szalay estimator (Landy & Szalay, 1993) is used for all auto-correlation measurements. Finally, in the case of the CCs, which use the Davis & Peebles (1983) estimator, only one random catalogue is needed. Hence, one can decide whether to use a random catalogue for the spectroscopic or KiDS data. We decided to use random catalogues for the spectroscopic data instead of the KiDS data in those cases, since most of the spectroscopic surveys provide well-established random catalogues that take into account and correct for a lot of systematic effects.
@@ 5.2 Fiducial analysis setup
We measured angular correlations in a single bin of fixed transverse physical separation between $0.5<r\leq 1.5\penalty 10000\ \mathrm{Mpc}$ in 32 linearly spaced redshift bins in the range $0.05\leq z<1.6$. [^11] These scales are larger than what was used in previous KiDS clustering redshift analyses. See Sect. 6.2.2 for a justification. For MICE2, the upper redshift limit is reduced to $z_{\rm max}=1.4$.
One of the main goals of any clustering-redshifts setup is to suppress and mitigate the redshift evolution of the galaxy bias of the calibration data and the target wide-field dataset. Following the notation of van den Busch et al. (2020), these biases can, under certain assumptions, be expressed in terms of the amplitudes of the angular autocorrelation functions of our spectroscopic reference sample, $w_{\rm ss}(z)$, and the wide-field photometric sample, $w_{\rm pp}(z)$. Then, the true (unknown) redshift distribution can be written as
@@
where $w_{\rm sp}(z)$ is the CC amplitude between the spectroscopic and photometric samples, $\Delta z$ is the bin width of the CC measurements, and $N_{\rm CC}(z)$ denotes our estimated $N(z)$ from the CC method.
In our fiducial analysis we chose to only correct for the calibration data bias; that is, we effectively measure the numerator of the right-hand side of Eq. (4). This is achieved by measuring the angular auto-correlation function of the calibration sample in the same redshift and distance bins as the CC function. This approach has been shown to successfully remove the redshift evolution of the galaxy bias of that sample. However, a similar treatment is not possible for the KiDS-Legacy source samples. The impact of the unknown galaxy bias of that sample data is partly mitigated by the fact that we bin the data tomographically (Schmidt et al., 2013), and it has been shown previously to be sufficiently small that we can neglect it in our following analysis (see e.g. Hildebrandt et al., 2021; van den Busch et al., 2020).
Finally, we produced a joint redshift estimate from all five calibration samples for which we measured the correlation functions independently. We obtained this final estimate by computing the inverse-variance weighted average of the bias-corrected correlation measurements of all reference samples. This produces an optimal redshift estimate that reflects the redshift-dependent relative contribution of each correlation measurement to the overall redshift estimate and accounts for the different statistical power each calibration sample has at any given redshift (Fig. 7). In general, the joint CCs were found to be dominated by BOSS, GAMA, and 2dFLenS at $z\lesssim 0.7$ and by DESI at high redshifts, whereas VIPERS had an overall small contribution due to its limited area and number density.
Figure 7: Example of the inverse-variance weighted combination of $C$s computed from the fourth tomographic bin of MICE2. The top panel shows the individual measurements from each calibration sample, the middle panel the weighted average, and the bottom panel the relative weight of each sample as a function of redshift.
@@ 5.3 Adaptation of signal-to-noise in MICE2
When comparing the S/N in the measured CCs on the between the data (from KiDS and KiDZ) and MICE2, we find that the noise level in the simulation is typically much smaller in most redshift bins, especially in the sixth tomographic bin (see top panel of Fig. 8). The reason for this difference is not entirely clear. Possible reasons could be differences in the selection functions applied to MICE2 for both the calibration and KiDS-Legacy data, as compared to the real data. There may also be a difference in the overall clustering amplitude found in the data and the simulations, especially on small scales. Additionally, the key difference between the data and MICE2 is that the mock photometry is perfectly uniform such that the effects of variable depth (Heydenreich et al., 2020) are not present in the simulation. Finally, there may be other systematic effects in the data for which our MICE2 data does not account, for example, related to fibre-collisions in the spectroscopic reference sample, for which DESI (for example) require special pair-weights (Bianchi et al., 2018) that we currently cannot integrate into our correlation estimator.
Since there was no clear explanation for this difference between data and simulation, we opted to adapt the measurements on the mocks by adding additional Gaussian noise to the values and inflating the error bars such that they match the data. We computed the right amount of noise required from the difference between the covariances of the inverse-variance-combined measurements of data and mock. Figure 8 shows a comparison of the original measurements in MICE2 and how the adapted version compares to the data measurements in the sixth tomographic bin.
Figure 8: One realisation of the CC measurements from MICE2 in the sixth tomographic bin after noise adaptation (Sect. 5.3) and the addition of intrinsic scatter (Sect 5.4). The top panel compares only the uncertainties of $N_{\rm CC}(z)$ from the data (grey) and MICE2 before (blue) and after adaptation (red). The dashed black line indicates the fitted intrinsic scatter from the measurements on the data ($f=0.14\pm 0.04$; see Eq. 6). The bottom panel shows the measured $N_{\rm CC}(z)$ from MICE2 (blue points) compared to the data (grey line). The red data points represent the MICE2 measurements after adapting the noise (up-scaling errors and perturbing values) and adding the intrinsic scatter (only perturbing valu-es) obtained from the data.